A 20-m long GC column was operated at 80 °C. A small peak for air had a retention time of 419 s and a base peak width of 14.8 s. A peak for heptane had a retention time of 536 s and a base peak width of 19.4 s. What is the chromatographic resolution between the air and heptane peaks?
Answer #1

Resolution (R) = 2(tR2 - tR1) / Wb1+ Wb2

          Where,   

                        tR1 = Retention time of air

                        tR2 = Retention time of heptane

                        Wb1 = Base peak width of air

                        Wb2 = Base peak width of heptane

Putting all the velues in the formula, we get

              R = 2(536-419) / 14.8 + 19.4

              R = 2(117) / 34.2

              R = 6.84 Answer